# New constraints on singlet scalar dark matter model with LZ, invisible Higgs decay and gamma-ray line observations

### v1

@@ Abstract
The singlet scalar dark matter (DM) model is a minimal extension of the Standard Model (SM). This model features only two free parameters: the singlet scalar mass $m_{S}$ and the quartic coupling $a_{2}$ between the singlet scalar and the SM Higgs doublet. Previous studies have suggested that only the resonant region remains viable under combined constraints. In this work, we revisit and refine the constraints using data from the direct detection experiment LZ, invisible Higgs decay measurements, and gamma-ray line observations conducted by Fermi-LAT and DAMPE. Our findings indicate that the latest results from LZ have significantly strengthened the constraints on the model parameters. We demonstrate that only a narrow parameter region remains viable, specifically $60.5\,\text{GeV}<m_{S}<62.5\,\text{GeV}$ and $1.7\times 10^{-4}<a_{2}<4.7\times 10^{-4}$. This suggests that a three-fold improvement in the current LZ results would thoroughly examine the parameter space below $1\,\rm{TeV}$ for the singlet scalar DM model.
@@ I Introduction
Many cosmological and astronomical observations provide compelling evidence for the existence of dark matter (DM), yet its nature remains a mystery. The Weakly Interacting Massive Particle (WIMP) is considered one of the most promising candidates for DM due to its ability to accurately reproduce the observed DM relic abundance through the so-called thermal relic mechanism Bergstrom:2000pn ; Bertone:2004pz ; Arcadi:2017kky . In recent years, indirect detection of DM through cosmic-ray searches has revealed several intriguing excesses, including the Galactic Center GeV gamma-ray excess (GCE) Hooper:2010mq , AMS-02 antiproton excess Cui:2016ppb ; Cuoco:2016eej , and electron/positron excess Chang:2008aa ; HESS:2009chc ; Fermi-LAT:2009yfs ; PAMELA:2008gwm ; Fermi-LAT:2011baq ; AMS:2013fma ; AMS:2014xys . These anomalies can be coherently interpreted as resulting from WIMP annihilation into SM final statesFeng:2013zca ; Zhou:2014lva ; Calore:2014xka ; Daylan:2014rsa ; Cui:2018klo ; Cholis:2019ejx .
The simplest model for a WIMP candidate is the Singlet Scalar Higgs portal model, which only introduces a new massive real scalar field $S$ that does not carry any charges under the Standard Model (SM) gauge groups Silveira:1985rk ; McDonald:1993ex ; Burgess:2000yq ; Davoudiasl:2004be ; OConnell:2006rsp . The discrete $Z_{2}$ symmetry is introduced to ensure the stability of the new scalar field. Under this symmetry, the SM Higgs doublet $H$ is $Z_{2}$-even while the new scalar $S$ is $Z_{2}$-odd. The symmetries of the Standard Model restrict the renormalizable interactions between $S$ and the SM fields to terms of the form $S^{2}H^{\dagger}H$ in the Lagrangian, as indicated by Eq.(1). This interaction allows $S$ to couple with the SM via the Higss field, resulting in various observable effects. Specifically, $S$ can annihilate into SM particles via the Higgs portal. This annihilation process is essential for the DM thermalization in the early Universe and can produce annihilation signals in the Milky Way today. Furthermore, DM-quark or gluon elastic scattering mediated by Higgs exchange could yield detectable signals in direct detection experiments. Finally, the SM Higgs boson has the potential to decay into a pair of $S$, which could be tested in collider experiments. Early related studies can be found in Refs. 1009.5377 ; 1306.4710 ; Feng:2014vea ; 1407.6588 ; 1609.03551 ; 1705.07931 ; 2101.02507 ; 2305.11937 .
@@ II Model Setup
For direct detection, we utilized the latest data from the LZ experiment to constrain the model parameters. The DM-nucleus spin-independent cross section for this model is given by Cheung:2014ppa
@@
The indirect detection constraints arise from the gamma-ray line spectrum. We focus on the DM annihilating into two photons, represented by the Feynman diagram in Fig.1. The cross section for this process is given by Cline:2013gha ; LHCHiggsCrossSectionWorkingGroup:2011wcg ; Gunion:1989we ; Djouadi:2005gi
@@ III Result
Following the analysis method in Ref. Feng:2014vea , we define $\Delta=(2m_{S}-m_{h})/m_{h}$ to focus on the resonant region, and the results are presented in Fig.3. The gamma-ray spectrum data effectively exclude the parameter space for masses slightly above $m_{h}/2$, but the exclusion capability is weaker for the region just below $m_{h}/2$. This is consistent with previous findings Feng:2014vea , though the improvement from the new gamma-ray line date is not substantial.
For DM masses slightly below $m_{h}/2$, the direct detection experiments LZ is particularly effective in excluding a substantial portion of the parameter space. The viable region is defined by the range $-10^{-3.4}<\Delta<-10^{-1.5}$, corresponding to $60.5\,\text{GeV}<m_{S}<62.5\,\text{GeV}$. Within this region, the coupling parameter is approximately $1.7\times 10^{-4}<a_{2}<4.7\times 10^{-4}$. If the LZ data were to improve by a factor of three, it could effectively access these parameter spaces. Moreover, if the DARWIN experiment achieves its expected precision, it could thoroughly probe this region.

### v2

@@ Abstract
The singlet scalar dark matter (DM) model is a minimal extension of the Standard Model (SM). This model features only two free parameters: the singlet scalar mass $m_{S}$ and the quartic coupling $a_{2}$ between the singlet scalar and the SM Higgs doublet. Previous studies have suggested that only the resonant region remains viable under combined constraints. In this work, we revisit and refine the constraints using data from the direct detection experiment LUX-ZEPLIN(LZ), invisible Higgs decay measurements, and gamma-ray line observations conducted by Fermi-LAT and DAMPE. Our findings indicate that the latest results from LZ have significantly strengthened the constraints on the model parameters. We demonstrate that only a narrow parameter region remains viable, specifically $60.5\,\text{GeV}<m_{S}<62.5\,\text{GeV}$ and $1.7\times 10^{-4}<a_{2}<4.7\times 10^{-4}$. This suggests that a three-fold improvement in the current LZ results would thoroughly examine the parameter space below $1\,\rm{TeV}$ for the singlet scalar DM model.
@@ I Introduction
Many cosmological and astronomical observations provide compelling evidence for the existence of dark matter (DM), yet its nature remains a mystery. The Weakly Interacting Massive Particle (WIMP) is considered one of the most promising candidates for DM due to its ability to accurately reproduce the observed DM relic abundance through the so-called thermal relic mechanism Bergstrom:2000pn ; Bertone:2004pz ; Arcadi:2017kky . In recent years, indirect detection of DM through cosmic-ray searches has revealed several intriguing excesses, including the Galactic Center GeV gamma-ray excess (GCE) Hooper:2010mq , AMS-02 antiproton excess Cui:2016ppb ; Cuoco:2016eej , and electron/positron excess Chang:2008aa ; HESS:2009chc ; Fermi-LAT:2009yfs ; PAMELA:2008gwm ; Fermi-LAT:2011baq ; AMS:2013fma ; AMS:2014xys . These anomalies can be coherently interpreted as resulting from WIMP annihilation into SM final statesFeng:2013zca ; Zhou:2014lva ; Calore:2014xka ; Daylan:2014rsa ; Ge:2017tkd ; Cui:2018klo ; Cholis:2019ejx ; Ge:2020tdh ; Abdughani:2021oit ; Fan:2022dck ; Chen:2024njd ; Fan:2024wvo ; Fan:2024rcr .
The simplest model for a WIMP candidate is the Singlet Scalar Higgs portal model, which only introduces a new massive real scalar field $S$ that does not carry any charges under the Standard Model (SM) gauge groups Silveira:1985rk ; McDonald:1993ex ; Burgess:2000yq ; Davoudiasl:2004be ; OConnell:2006rsp . The discrete $Z_{2}$ symmetry is introduced to ensure the stability of the new scalar field. Under this symmetry, the SM Higgs doublet $H$ is $Z_{2}$-even while the new scalar $S$ is $Z_{2}$-odd. The symmetries of the Standard Model restrict the renormalizable interactions between $S$ and the SM fields to terms of the form $S^{2}H^{\dagger}H$ in the Lagrangian, as indicated by Eq.(1). This interaction allows $S$ to couple with the SM via the Higss field, resulting in various observable effects. Specifically, $S$ can annihilate into SM particles via the Higgs portal. This annihilation process is essential for the DM thermalization in the early Universe and can produce annihilation signals in the Milky Way today. Furthermore, DM-quark or gluon elastic scattering mediated by Higgs exchange could yield detectable signals in direct detection experiments. Finally, the SM Higgs boson has the potential to decay into a pair of $S$, which could be tested in collider experiments. Early related studies can be found in Refs. Profumo:2010kp ; 1306.4710 ; Feng:2014vea ; 1407.6588 ; 1609.03551 ; 1705.07931 ; 2011.13225 ; Ren:2014mta ; 2101.02507 ; 2305.11937 ; 2406.01705 .
@@ II Model Setup
For direct detection, we utilized the latest data from the LZ experiment to constrain the model parameters. The DM-nucleus spin-independent cross section for this model is given by Cheung:2012xb
@@
The indirect detection constraints arise from the gamma-ray line spectrum. We focus on the DM annihilating into two photons, represented by the Feynman diagram in Fig.1. The cross section for this process is given by 1306.4710 ; LHCHiggsCrossSectionWorkingGroup:2011wcg ; Gunion:1989we ; Djouadi:2005gi
@@ III Result
In the Higgs resonance region, i.e., $m_{S}\simeq m_{h}/2$, the allowed range of the coupling coefficient $a_{2}$ that satisfies the correct DM relic density exhibits a sharp drop. As show in Fig. 1, in this model, DM annihilation occurs via the $s$-channel Higss exchange. According to Eq. (5), when $s=(2m_{S})^{2}\simeq m_{h}^{2}$, the denominator term reaches its minimum, significantly enhancing the DM annihilation cross-section. Consequently, this resonance mechanism enables even a small coupling ($a_{2}\sim 10^{-4}$) to drive sufficient annihilation in the early universe, ensuring consistency with the relic density constraint.
Notably, the Higss resonance annihilation can generate a distinctive gamma-ray line signal, to which DM indirect detection experiments are highly sensitive. To better illustrate the latest constraints from DAMPE and Fermi-LAT gamma-ray line searches on the resonance parameters, we define $\Delta=(2m_{S}-m_{h})/m_{h}$ following Ref. Feng:2014vea to focus on the resonant region, with the corresponding results presented in Fig. 3. The gamma-ray spectrum data strongly exclude the parameter space for masses slightly above $m_{h}/2$, but the exclusion power is weaker for the region just below $m_{h}/2$. This is because, during the freeze-out stage in the early universe, DM $S$ carries non-negligible kinetic energy, allowing the resonance condition $s=m_{h}^{2}$ to be met even for $m_{S}<m_{h}/2$. As a result, the relic density constraint favors $m_{S}$ slightly below $m_{h}/2$. However, in the present-day Galactic center, where DM is nearly non-relativistic, the resonance condition holds at $m_{S}\approx m_{h}/2$, making indirect detection constraints more effective for masses above the resonance.
For DM masses slightly below $m_{h}/2$, the direct detection experiments LZ LZCollaboration:2024lux is particularly effective in excluding a substantial portion of the parameter space. The viable region is defined by the range $-10^{-3.4}<\Delta<-10^{-1.5}$, corresponding to $60.5\,\text{GeV}<m_{S}<62.5\,\text{GeV}$. Within this region, the coupling parameter is approximately $1.7\times 10^{-4}<a_{2}<4.7\times 10^{-4}$. If the LZ data were to improve by a factor of three, it could effectively access these parameter spaces. Moreover, if the DARWIN DARWIN:2016hyl experiment achieves its expected precision, it could thoroughly probe this region.